OpenStudy (anonymous):

There are students in a speech contest. Yesterday, of them gave their speeches. Today, of the remaining students gave their speeches. How many students still haven't given their speeches?

OpenStudy (dan815):

You are missing the fractions or something there before the " of them"

OpenStudy (anonymous):

There are 72 students in a speech contest. Yesterday, 1/4 of them gave their speeches. Today 2/3 of the remaining students gave their speeches. How many students still haven't given their speech? (1/4)*72 = 72/4 = 18 students gave their speech yesterday So there are 72 - 18 = 54 students who need to give their speech 2/3 of those 54 students gave their speech today, so (2/3)*54 = (2*54)/3 = 108/3 = 36 students gave their speech today

OpenStudy (goformit100):

Warm Welcome To Open Study